2. The Accounting Cycle – End of the period (7pts): Following are the accounts and balances that appear in the unadjusted trial balance for Just Like Fred Astaire (aka ‘Fred’) dance studio as of 9/30/2019. Assume all accounts have their normal debit or credit balance:

Account:

Amount:

Account:

Amount:

Cash

$18,944

Retained Earnings

$41,392

Accounts Receivable (A/R)

$15,088

Service Revenue

$89,430

Prepaid Rent

$10,100

Advertising Expense

$12,550

Equipment

$37,920

Wages Expense

$48,000

Accumulated Depreciation-Equipment

$16,195

Utilities Expense

$2,775

Accounts Payable (A/P)

$5,901

Depreciation Expense

$0

Wages Payable

$0

Rent Expense

$0

Deferred Revenue

$240

Cleaning Expense

$9,040

Common Stock

$1,259

The following additional information was gathered at September 30, 2019, the end of Just Like Fred Astaire’s quarterly accounting period. Assume Fred uses a quarterly accounting period and all adjustments and adjusting entries are made only at the end of each quarterly accounting period (i.e. AJEs are only made on March 31, June 30, September 30 and December 31st).   No adjusting entries have been made yet for the third quarter, ending September 30th, 2019 related to the below items:

  1. On August 1, 2019, Fred prepaid $8,700 for the next six months’ rent, covering the period of August 1, 2019 – May 31, 2020. This was recorded as Prepaid Rent. Fred last prepaid rent on February 1, 2019 in the amount of $8,400, which covered the six-month period of February 1, 2019 – July 31, 2019.
  2. Fred charges customers $60 for each dance class. On September 15th, Daniel registered for three group classes with Fred. One was provided in September, on September 25th. The other two lessons will be provided in the following accounting period, on October 2nd and October 9th. Daniel will pay for all three classes on October 9th. Fred has not yet billed Daniel for the September class, and no entry related to the September class already provided has been recorded yet.
  3. Fred purchased a truck for $37,920 on February 1, 2016, which was recorded as Equipment. The truck has $0 salvage value, and has an estimated useful life of 8 years. Fred uses the straight-line depreciation method to allocate the cost of the truck evenly over its eight-year useful life.
  4. Fred pays its workers a total of $8,000 every 2-week pay period. Each 2-week pay period includes exactly 10 work days, i.e. $800 per work day. The last cash payment to workers was made on Friday September 20th for the pay period of September 7th – September 20th. The next cash payment of $8,000 will be made on October 4th and covers the pay period of September 21st – October 4th, which includes 6 work days left in September and 4 work days in October.
  5. Fred charges customers $60 for each dance class. On July 28th, Grace signed up for four separate group dance classes, one which was provided in August, two which were provided in September, and one which will be provided in October. Grace prepaid Fred $240 in advance for all four classes on July 28th which was recorded as Deferred Revenue.

PART A: Record Adjusting Journal Entries (AJEs) – 5pts: Record the five AJEs related to the information above that Fred would make on 9/30/19. Make sure to use proper formatting for all entries, and to include the date of each entry and a brief description of each entry. You must show all of your work on calculations to receive full credit.

PART B: Calculate adjusted Net Income for 2017 – 1pts: Calculate the amount of Net Income that Fred would report on its income statement for the period ended September 30, 2019.   You must show all of your work on calculations to receive full credit.

PART C: Calculating Net Book Value – 1pt: Calculate the net book value of the equipment that would be reported on the classified Balance Sheet as of 9/30/2017. You must show all of your work on calculations to receive full credit.
Answer #1

PART A - The journal entries to be booked will be:

Date Credit ($) Debit ($) 4,300 Sep 30 15,000 Sep 30 60 60 Sep 30 1,185 Account Titles & Description Rent Expense Prepaid Ren

PART B - Net Income = $7,020

PART C = $20,540

Workings

Date Sep 30 Debit ($) Credit ($) workings & logic 4,300 =(8400*1/6)+(8700*2/6) 15,000 Sep 30 60 revenue booked only for class

Net Income = Revenues- Expenses = (89430+60+180) - (12550+48000+2775+9040+4300+1185+4800) = 89670-82650 = $7,020

Net book value of asset = 37920-(16195+1185) = 37920-17380 = $20,540

Please hit LIKE button if this helped. For any further explanation, please put your query in comment, will get back to you.